# Light-A-Video: Training-free Video Relighting via Progressive Light Fusion

### v1

@@ 5 Experiments
Figure 5: Qualitative comparison of baseline methods. Given a source video and guidance text prompt, we compare the performance of Light-A-Video with other video editing methods. VDM used: AnimateDiff (Left), CogVideoX (Right).
In this section, we provide comprehensive qualitative and quantitative evaluations to highlight the efficacy and generality of our method for zero-shot video illumination control. Furthermore, we conduct ablation studies to analyze the contributions of critical model components and provide additional results on text-conditioned video illumination modifying with background generation.
@@ 5.1 Experimental Details
Baselines. In the absence of established video relighting methods, we adopt the state-of-the-art image relighting technique to perform frame-by-frame relighting on videos as a baseline. To verify the temporal smoothing effect of illumination using a VDM, we add 20% and 60% noise to the results obtained from IC-Light [63]. Subsequently, we apply SDEdit [31] on VDM for denoising, resulting in two sets of results named IC-Light + SDEdit-0.2 and IC-Light + SDEdit-0.6, respectively. Finally, AnyV2V [26] is utilized as another baseline to verify whether the temporal content migration method can be effectively applied to video relighting tasks. This baseline names IC-Light + AnyV2V, where the first frame of the IC-Light result and the source video are used as conditional inputs.
@@
Evaluation metrics. We report three widely adopted metrics for quantitative evaluation. First, to evaluate the temporal consistency of the generated video, the average CLIP [39] score across consecutive video frames is utilized. Second, RAFT [50] is used to estimate the optical flow for each baseline video, and assess the motion preservation of each method by calculating the optical flow similarity with the source video. Third, to evaluate the quality of image relighting for each method, a video test dataset is collected, and the FID [43] score is calculated between each method’s results And the frame-by-frame IC-Light results, serving as the relight quality evaluation metric.
@@
Datasets. We constructed a video test dataset consisting of 73 videos. The majority of these videos were selected from the DAVIS [37] public dataset, which contains a diverse collection of semantically rich videos with pronounced motion. Additionally, some videos were collected from Pixabay [36], featuring high-quality videos with significant motion. All quantitative metrics are evaluated on our collected dataset. For each video, two lighting prompts are applied, and three lighting directions are randomly chosen.
@@
Implementation details. If not specified, the default models employed for image relighting and VDM in the subsequent experiments are IC-Light [63] and AnimateDiff [13] motion-adapter-v3, respectively. In the IC-Light model, the lighting conditions $c$ for image relighting are derived from two components. First, a text prompt that describes the characteristics of the light source (e.g., neon light, sunshine, etc.). Second, a lighting map is utilized to represent the light intensity across the scene. This lighting map is then encoded by a VAE and serves as the initial latent for the denoising process. During the inference stage, the source video is added with 50% noise. Then, the VDM employs a denoising process with $T_{m}=25$ steps to progressively fusion the relight target. For the parameters in our Light-A-Video model, the $\gamma=0.5$ in the CLA module is used to balance the original attention hidden state and the cross-frame averaged hidden state. In our PLF strategy, the fusion weight $\lambda_{t}$ is decreases as denoising progresses, so we set $\lambda=1-t/T_{m}$. All experiments were conducted on a single A100 GPU, with a memory consumption of approximately 23GB for a 16-frame relighted video generation.
@@
Figure 7: Ablation Study. We present the video relighted with background generation results by removing the CLA module or the PLF strategy.
5.2 Qualitative Results
As shown in Fig. 5, the frame-by-frame IC-Light method ensures single-frame quality. However, the absence of consistency design and VDM temporal priors results in significant flickering of the light source and overall appearance. By introducing VDM priors, IC-Light + SDEdit-0.2 maintains content consistent with the source video, but still exhibits noticeable relight appearance jitter. IC-Light + SDEdit-0.6 further enhances temporal smoothness, yet object identity shifts occur. AnyV2V can transfer the appearance of the first relight frame to subsequent frames, but this type of temporal content migration method inherently performs only pixel-level migration without perceiving the given light source, leading to unreasonable changes in object illumination. In contrast, Light-A-Video achieves high-quality video relighting, demonstrating strong temporal consistency and high fidelity to the light source.
5.3 Video Relighting with Background Generation
As depicted in Fig. 6, Light-A-Video not only supports the direct relighting of the input source video but also accepts a video foreground sequence and a text prompt provided by users as input, generating a corresponding video background and illumination that aligned with the prompt descriptions. Specifically, we first process the input foreground sequence with IC-Light for frame-by-frame relighting, while completely noising the background to serve as the initialization video for VDM. From step $T$ to $T_{m}$, the background is gradually denoised, leveraging VDM’s inpainting capability to generate the background. Then, from step $T_{m}$ to 0, CLA and PLF are introduced to achieve a temporally consistent relighting appearance of the video. The results in Fig. 6 demonstrate that our pipeline can achieve high-quality video relighting results with consistent background generation.
@@ 5.4 Quantitative Evaluation
The quantitative comparison of our method with various baselines is presented in Tab. 1. As shown, we introduced an additional baseline SDEdit-0.6, which denoises the source video with 60% noise. SDEdit-0.6 results in a high temporal Clip Score demonstrating relying solely on the relight prompt is insufficient for generating videos with the corresponding lighting characteristics using VDM. When the frame-by-frame IC-Light results are used as the initialization video and then noised, IC-Light + SDEdit-0.6 exhibits an increased FID score, indicating a trade-off with temporal stability. Similarly, IC-Light + SDEdit-0.2, which incorporates 20% noise, maintains a high consistency in lighting effects with frame-by-frame IC-Light, but it also suffers from unacceptable temporal flickering. For AnyV2V, a temporal content migration method, the first frame’s appearance is consistent with the IC-Light results. Yet, its inability to perceive the light source, coupled with the model’s inherent quality degradation in subsequent frames, results in a low motion preservation score. In contrast, our Light-A-Video method achieves a low FID score while maintaining a high level of temporal consistency, demonstrates the effectiveness in both relighted image quality and temporal stability.
Evaluation Metric (a) Image Quality (b) Temporal Consistency
FID Score($\downarrow$) CLIP Score($\uparrow$) Motion Preservation($\downarrow$)
IC-Light [63] / 0.9040 5.969
SDEdit-0.6 [31] 64.03 0.9677 4.204
IC-Light + SDEdit-0.2 13.79 0.9199 5.959
IC-Light + SDEdit-0.6 62.61 0.9483 7.544
IC-Light + AnyV2V [26] 32.73 0.9436 8.854
Light-A-Video (Ours) 29.63 0.9655 2.493
Table 1: Quantitative comparison of baseline methods. The best result is highlighted in bold, while the second-best result is underlined.
@@ 5.5 Ablation Study
We conduct an ablation study to understand the importance of our CLA and PLF modules. As shown in Fig. 7, for the video relighting with background generation task, frame-by-frame IC-Light ensures high single-frame relighting quality but lacks temporal consistency control, leading to lighting source inconsistency and relighted appearance inconsistency. With the help of CLA, cross-frame information exchange stabilizes the generation of background lighting sources. Furthermore, by introducing VDM motion priors and using PLF’s strategy to progressively fuse relight targets into the original denoising target, Light-A-Video achieves temporally smooth relighting of the video. The overall video quality is also significantly improved with the help of VDM priors.
5.6 Limitation and Future Work.
Despite achieving impressive results, the performance of our training-free method is limited by the capabilities of both the image-relighting model and the VDM. While Light-A-Video excels at ensuring stable lighting and temporal consistency, the CLA module, which stabilizes the background lighting, is less effective for modeling dynamic lighting changes. To address this limitation, future work will focus on developing methods that can handle dynamic lighting conditions more effectively.
@@ 6 Conclusion
In summary, our paper introduces Light-A-Video, a training-free method that leverages the capabilities of state-of-the-art image relighting models to achieve temporally consistent video relighting. By integrating the Consistent Light Attention to stabilize lighting source generation and employ the Progressive Light Fusion strategy for smooth appearance transitions. Light-A-Video significantly enhances the temporal coherence of relighted videos while preserving the high-quality relighting of images. Experiments demonstrate the effective application of advanced image-relighting techniques in both video sequence relighting and foreground relighting with background generation.

### v2

@@ 5 Experiments
Figure 6: Qualitative comparison of baseline methods. Given a source video and guidance text prompt, Light-A-Video achieves high temporal consistency and fidelity to the light condition, outperforming other methods in avoiding flickering, jitter, and identity shifts. VDM used: AnimateDiff (Left), CogVideoX (Right).
Evaluation Metric (a) Relighted Image Quality (b) Temporal Consistency (c) User Preference
FID Score ($\downarrow$) CLIP Score ($\uparrow$) Motion Preservation ($\downarrow$) LPA ($\uparrow$) VS ($\uparrow$) IP ($\uparrow$)
IC-Light [62] / 0.9040 5.969 3.160 2.128 3.068
IC-Light + SDEdit-0.2 13.79 0.9199 5.959 2.850 2.752 3.014
IC-Light + SDEdit-0.6 62.61 0.9483 7.544 2.472 3.318 2.488
IC-Light + AnyV2V [25] 32.73 0.9436 8.854 2.766 3.300 2.774
Light-A-Video (Ours) 29.63 0.9667 1.833 3.752 3.502 3.656
Table 1: Quantitative comparison of baseline methods. We achieves better results in relighted image quality and temporal stability.
@@ 5.1 Experimental Details
Baselines. Given the lack of established video relighting methods, we adopt the state-of-the-art image relighting technique to perform frame-by-frame relighting on videos as a baseline. To verify the temporal smoothing effect of illumination using a VDM, we construct two comparative methods by applying SDEdit [30] to the per-frame results of IC-Light [62]. These two methods are named IC-Light + SDEdit-0.2 and IC-Light + SDEdit-0.6, corresponding to noise levels of 20% and 60%. Finally, IC-Light + AnyV2V [25] serves as another baseline. Specifically, IC-Light relights the first frame, and AnyV2V propagates the appearance information from the first frame to all subsequent frames, preserving the content of the source video.
@@
Evaluation metrics. Three widely adopted metrics are reported for quantitative evaluation. First, the temporal consistency of the generated video is assessed using the average CLIP [38] score across consecutive video frames. Second, the optical flow for each baseline video is estimated using RAFT [49], and the motion preservation score of each method is assessed by calculating the optical flow discrepancy with the source video. Third, to evaluate the quality of relighted image, a video test dataset is collected. The FID [42] score is then calculated between the results of each method and the frame-by-frame IC-Light results, serving as the metric for relight quality evaluation. Finally, 52 volunteers are invited to conduct a user study across three aspects: Lighting Prompt Alignment (alignment between video content and lighting prompt), Video Smoothness (temporal consistency of the relighted video), and ID-Preservation (consistency of the object’s identity and albedo before and after relighting). The volunteers rank the results of five methods, and the average user ranking is used as a preference metric.
@@
Datasets. We constructed a video test dataset consisting of 73 videos. The majority of these videos are selected from the DAVIS [36] public dataset, which contains a diverse collection of semantically rich videos with pronounced motion. Additionally, some videos are collected from Pixabay [35], featuring high-quality videos with significant motion. All quantitative metrics are evaluated on our collected dataset. For each video, two lighting prompts are applied, and three lighting directions are randomly chosen.
@@
Implementation details. Unless otherwise specified, the default models employed for image relighting and VDM in the subsequent experiments are IC-Light [62] and AnimateDiff [13] Motion-Adapter-v3, respectively. In the IC-Light model, the lighting conditions $c$ for image relighting are derived from two components: First, a text prompt that describes the characteristics of the light source (e.g., neon light, sunshine, etc.). Second, a lighting map is utilized to represent the light intensity across the scene. This lighting map is then encoded by a VAE and serves as the initial latent for the denoising process. During the inference stage, the source video is added with 50% noise. Subsequently, the VDM employs a denoising process with $T_{m}=25$ steps to progressively fuse the relight target. For the parameters in the pipeline, $\gamma=0.5$ in the CLA module is used to balance the original attention feature and the cross-frame averaged feature. In the PLF strategy, the fusion weight $\lambda_{t}$ decreases as denoising progresses, and we set $\lambda_{t}=1-t/T_{m}$.
5.2 Qualitative Results
As depicted in Fig. 6, the frame-by-frame IC-Light method ensures high single-frame quality. However, the lack of consistency design and VDM temporal priors leads to significant flickering of the light source and overall appearance. By introducing VDM priors, IC-Light + SDEdit-0.2 maintains content consistent with the source video, but still exhibits noticeable relight appearance jitter. IC-Light + SDEdit-0.6 further enhances temporal smoothness, yet object identity shifts occur. AnyV2V transfers the appearance of the first relight frame to subsequent frames, but this pixel-level migration method, lacking perception of the given light source, results in unreasonable illumination changes. In contrast, Light-A-Video achieves high-quality video relighting, demonstrating strong temporal consistency and high fidelity to the light source.
5.3 Video Relighting with Background Generation
As depicted in Fig. 7, Light-A-Video can accept a video foreground sequence and a user-provided text prompt as input, generating a corresponding video background and illumination that aligns with the prompt descriptions. Specifically, the input foreground sequence is initially processed with IC-Light for frame-by-frame relighting, while the background is entirely noised to serve as the initialization latent for the VDM. From step $T$ to $T_{m}$, the background is progressively denoised, leveraging the VDM’s inpainting capability to generate the background. Subsequently, from step $T_{m}$ to 0, the CLA module and PLF strategy are employed to achieve a temporally consistent relighting appearance of the video. These results illustrate that our pipeline can produce high-quality video relighting results with consistent background generation.  [^1] More examples and ablation experiments are provided in the supplementary material.
@@ 5.4 Quantitative Evaluation
The quantitative comparison of our method with various baselines is presented in Tab. 1. Given the addition of only 20% noise, IC-Light + SDEdit-0.2 exhibits performance in video relighting that is nearly identical to IC-Light, resulting in significant temporal flickering in both methods. IC-Light + SDEdit-0.6 provides enhanced temporal consistency but suffers from object identity shifts due to the introduction of excessive noise. For AnyV2V, the appearance of the first frame aligns well with the IC-Light results. However, its inability to perceive the light source, combined with inherent quality degradation in subsequent frames, leads to a low motion preservation score. In contrast, Light-A-Video achieves a low FID score while maintaining high temporal consistency, demonstrating its effectiveness in both relighted image quality and temporal stability.
Figure 8: Ablation Study. Results of video relighting with the CLA module or the PLF strategy removed.
@@ 5.5 Ablation Study
An ablation study is conducted to assess the importance of our CLA and PLF modules. As illustrated in Fig. 8, for the video relighting task involving background generation, frame-by-frame IC-Light provides high-quality single-frame relighting but lacks control over temporal consistency. This results in inconsistencies in lighting sources and relighted appearances across frames. The CLA module enables cross-frame information exchange, which stabilizes the generation of background lighting sources. Additionally, by introducing VDM motion priors and employing PLF’s strategy for progressive fusion of relight targets into the original denoising target, Light-A-Video ensures temporally smooth relighting. The overall video quality is also significantly improved with the aid of VDM priors.
5.6 Limitation and Future Work
Despite the impressive results achieved by our training-free method, its performance is inherently constrained by the capabilities of the underlying image-relighting model and the VDM. While Light-A-Video demonstrates remarkable proficiency in ensuring stable lighting and temporal consistency, the CLA module, which is designed to stabilize background lighting, exhibits limitations when it comes to modeling dynamic lighting changes. To address this limitation, future work will focus on developing novel methods that can more effectively handle dynamic lighting conditions.
@@ 6 Conclusion
In summary, this paper introduces Light-A-Video, a training-free method that utilizes state-of-the-art image relighting models to achieve temporally consistent video relighting. By incorporating the Consistent Light Attention (CLA) module to stabilize lighting source generation and employing the Progressive Light Fusion (PLF) strategy for smooth appearance transitions, Light-A-Video significantly enhances the temporal coherence of relighted videos while preserving the high-quality relighting of individual frames.
Supplementary Material
Figure 9: Evolution of $\lambda_{t}$ over time steps $t$ for different PLF strategies. $\lambda_{t}$ determines the proportion of the relight target mixed into the fusion target.
Appendix A Comprehensive Ablation Studies
In this section, we conduct comprehensive ablation studies to explore the effects of the hyper-parameter $\gamma$ of the Consistent Light Attention (CLA) and various Progressive Light Fusioin (PLF) strategies on the quality of relighted video generation. Specifically, the values of $\gamma$ are uniformly sampled within the range of $[0,1]$, where a larger $\gamma$ indicates a higher proportion of the cross-frame averaged feature in the CLA. Notably, when $\gamma=0$, it corresponds to the vanilla IC-Light with standard self-attention. For the PLF strategy, the parameter $\lambda_{t}$ determines the proportion of the relight target mixed into the fusion target at each step. Several different PLF strategies are also proposed, with $\lambda_{t}$ defined as:
$\lambda_{t}=1-\left(\frac{t}{T_{m}}\right)^{k}$ (14)
Here, $T_{m}=25$ denotes the total number of noise-adding steps for the source video, and different values of $k$ indicate different rates of decay for $\lambda_{t}$ over time. $\lambda_{t}\equiv 1$ means directly replacing the fusion target with the relight target for all steps. Fig. 9 illustrates the curves of $\lambda_{t}$ as it varies with time step $t$.
A quantitative comparison of various settings is provided in Fig. 10, where the three evaluation metrics (FID, Temporal Clip score, and Motion Preservation score) introduced in the main text are employed to evaluate the per-frame image quality and temporal consistency of the relighted video generated by our Light-A-Video method. Specifically, Fig. 10 (a) depicts the variation of the FID score with different values of the trade-off parameter $\gamma$. An excessively large $\gamma$ results in a significant degradation of the overall relighting image quality. This is attributed to the overemphasis on the cross-frame averaged feature in the CLA module, which leads to temporal over-smoothing and diminishes the lighting specificity, thereby negatively impacting the relighting effect. However, when $\gamma$ is chosen appropriately (between 0.2 and 0.5), the FID score remains stable and can even be enhanced, especially when employing PLF strategies with $k=1$ or $k=0.5$.
The temporal consistency evaluation, as depicted in Fig. 10 (b), demonstrates a steady increase in the Temporal Clip score with the rise of the parameter $\gamma$. This trend underscores the remarkable efficacy of the CLA module in augmenting the temporal consistency of the relighted video. These results reflect that the CLA module is highly effective in enhancing the temporal consistency of the relighted video. In a parallel vein, the Motion Preservation score serves as an indicator of motion consistency with the source video. Specifically, when the value of $\gamma$ is selected within the range of $0.2$ to $0.5$, the relighted video can achieve a high degree of motion consistency with the original video.
It is worth noting that, as evidenced by the three figures, employing a constant $\lambda_{t}\equiv 1$ significantly underperforms the method of progressively decreasing $\lambda_{t}$ in PLF, both in terms of relight image quality and temporal consistency. Although a constant $\lambda_{t}$ yields a higher Temporal Clip score when $\gamma>0.5$, the overall motion deviates substantially from the source video, resulting in an unacceptable motion preservation effect. These results effectively demonstrate the efficacy of our PLF strategy. The explanation for this observation is twofold:
  • Compared to a dynamically mixed target, a constant target with rich additional illumination information in the denoising process is more likely to deviate from the sampling trajectory of the Video Diffusion Model (VDM). When this deviation exceeds the refinement capability of the VDM, it perturbs the motion priors, consequently leading to visible temporal jitter.
  • Repeatedly injecting constant relight appearance across multiple iterations is analogous to cyclically relighting the same image using the image relight model. This process causes the input distribution to progressively diverge from the training distribution of the image relight model, ultimately degrading the quality of the relighted images.
Figure 10: The relative effectiveness of different PLF strategy on Light-A-Video performance. (a) FID scores, (b) Temporal CLIP scores, and (c) Motion Preservation scores are shown for four strategies: PLF with constant $\lambda$ ($\lambda_{t}\equiv 1$), and PLF with $k=0.5,1,2$. Lower FID/Motion Preservation scores and higher Temporal Clip scores indicate better performance.
Appendix B Additional Results
In this section, we present additional qualitative results. In Fig. [11-12], we show examples of foreground sequences relighting with background generation on AnimateDiff. In Fig. [13-14], we showcase the application of Light-A-Video directly to the video relighting task. And finally, as illustrated in Fig. 15, we present the video relighting results on DiT-based video models, such as CogVideoX.
Figure 11: More results of Light-A-Video in foreground sequences relighting with background generation.
Figure 12: More results of Light-A-Video in foreground sequences relighting with background generation.
Figure 13: More results of Light-A-Video in video sequences relighting.
Figure 14: More results of Light-A-Video in video sequences relighting.
Figure 15: More results of Light-A-Video in video sequences relighting on CogVideoX.